OpenStudy (anonymous):

Stephanie made 36 muffins. She made 24 apple muffins and the rest were banana. What is the ratio of apple muffins to banana muffins? A. 2 to 3 B. 24 to 36 C. 12 to 36 D. 2 to 1

OpenStudy (anonymous):

apple : banana

OpenStudy (anonymous):

\[\frac{ 24 }{ 36-24 } = \frac{ ? }{ ? }\]

OpenStudy (bookworm14):

24 apple so 36 - 24= #of banana muffins what does that give you?

OpenStudy (anonymous):

12

OpenStudy (bookworm14):

so it is 24 apple: 12 banana this makes the fraction \[\frac{ 24 }{ 12 }\] now reduce it :)

OpenStudy (anonymous):

is it d

OpenStudy (bookworm14):

\(\huge \color{green}{Correct~!}\color{blue}{\ddot\smile\heartsuit}\)

OpenStudy (anonymous):

thx

OpenStudy (bookworm14):

you divide the top ant bottom by 12 getting you 2:1 Your welcome!